OpenStudy (anonymous):

a student scores 71 out of 100 on a test. If the maximum score on the next test is also 100 points, what score does the student need to maintain at least an average of 79?

terenzreignz (terenzreignz):

You know how to get the average?

OpenStudy (anonymous):

no,please help

terenzreignz (terenzreignz):

The average of two numbers, a and b is given by \[\huge \frac{a+b}{2}\]

terenzreignz (terenzreignz):

So, we can let a = 71, as in his first test... \[\huge \frac{71+b}{2}\]

terenzreignz (terenzreignz):

And this average has to be at least 79, so... \[\huge \frac{71+b}{2}\ \ge \ 79\] Now you just have to solve for b :) Good luck :)

OpenStudy (anonymous):

huh?